Circular No. 6220

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) A. Fitzsimmons and M. Cartwright, Queen’s University of Belfast, have

observed the inner jet (IAUC 6216) in R-band CCD images with the 1-m Jacobus Kapteyn Telescope on La Palma; the jet appears nearly linear out to 3′′ from the nucleus, then turns sharply north. They see no significant change in p.a. (measured from the center of light to the turn) during Aug. 24.90 and 28.90 UT (305^{o}) and Aug. 25.93, 26.92, and 27.96 (315^{o}).

M. Kidger, M. Serra, and R. Casas, Instituto de Astrofisica de Canarias, report monitoring of the spiral-arm feature on Aug. 25 (2.5 hr in R), 27 (2.0 hr in R), and 28 (2.25 hr in B) with the 0.82-m IAC-80 Telescope at Teide Observatory: “The position angle (280^{o}) was as reported on IAUC 6216. No evidence is found for significant rotation of the base of the jet; these and observations on Aug. 31 show that the base oscillates by a maximum of 20^{o}–30^{o}, suggesting that the jet emanates from a point close to the pole of the nucleus (within 10^{o}–15^{o} latitude) and that the comet is traveling very close to pole-first. The arm extends considerably farther on Aug. 28 than on Aug. 25, wrapping round to p.a. 0^{o} on Aug. 25, 15^{o} on Aug. 27, and 25^{o} on Aug. 28. Significant changes were seen in the jet structure between Aug. 28.88 and 28.97, the distance between the nucleus and the end of the jet increasing from ∼ 8′′ to ∼ 10′′, while the p.a. of the end of the jet remained constant. On Aug. 31, the jet was much fainter (and, where it emerges from the nucleus, more tightly collimated).”

R. M. West, European Southern Observatory (ESO), writes: “E. Molinari (Osservatorio Astronomico di Brera) obtained 5-min exposures with the Danish 1.5-m telescope (+ multi-mode DFOSC) at La Silla on Aug. 31.06 (Gunn i filter; seeing 0′′.8) and Sept. 1.06 UT (Gunn r filter; 1′′.0);

see http://www.eso.org/comet-hale-bopp/comet-hale-bopp.html. On

both frames, the northwestern jet is well resolved and, following image processing of these frames by H.-U. Kaeufl (ESO-Garching), its structure may be described as follows: On the i frame, the jet emerges at 3′′ in p.a. 295^{o} relative to the point of maximum intensity. From there, it curves toward the north to a point ∼ 7′′ in p.a. 315^{o}, where it bends rather sharply ∼ 90^{o} toward the northeast, widens to FWHM ∼ 8′′, and continues in direction p.a. ∼ 40^{o} over a distance of ∼ 10′′. The spatial structure may be helical. A comparison of the two frames reveals no significant differences (on the r frame, the jet emerges at 3′′.5 in p.a. 293^{o}) and these images resemble those from Hawaii and Teide, now also displayed on the WWW. Thus, no obvious rotational effects can be perceived from these data.”

1995 September 1 Daniel W. E. Green